the mass of a particle is 10/7 kg and it is moving in x direction initial position x=0 m and velocity =1 power =2 W find the velocity at x=10 m power being 4 W

Grade:11

1 Answers

Thill
6 years ago
Power=work/time                  using these, w/t=p, 2=w/t, 2t=w, 2t=10/7*0, t=0, initially
work=mass*distance             in the same way, 4t=10/7*10, t=25/7.
distance=velocity*time          v=d/t=10/25/7=70/25=2.8